An angle measures 68° more than the measure of its supplementary angle. What is the measure of each angle
Answer by greenestamps(13258)



Formally....

Let x be the measure of the angle
Then 180-x is the measure of its supplement

The angle measure is 68 degrees more than the measure of the supplement:

x = (180-x)+68
x = 248-x
2x = 248
x = 124

ANSWERS:
the angle: x = 124 degrees
the supplement: 180-x = 180-124 = 56 degrees

Informally....

The sum of the measures of an angle and its supplement is 180 degrees.

That means the measures of the two angles differ from 90 degrees by equal amounts.

Since in this problem the difference between the measurements of the two angles is 68 degrees, the measures of the two angles differ from 90 degrees by 68/2 = 34 degrees.

So the measures of the two angles are 90+34 = 124 degrees and 90-34 = 56 degrees.

ANSWERS: 124 degrees and 56 degrees